# Leveraging Self-Supervised Learning for Speaker Diarization

### v2

@@ III-B2 Inference
Pyannote3 (baseline) SincNet 21.3 13.7 23.7 13.8 7.5 14.5
Proposed Fbank 19.8 12.2 21.0 13.1 6.6 12.4
WavLM-frozen 17.0 11.6 18.7 10.9 5.9 10.8
WavLM-updated 15.4 11.5 16.0 9.9 5.7 8.2
@@ III-C1 Overall comparison
It is clear that our EEND part has better performance than the Pyannote baseline. Even with the traditional filterbank features, our model significantly outperforms the Pyannote model across the three datasets, suggesting that the Conformer is a better choice for the end-to-end module. By utilizing WavLM, better performance can be obtained, especially for AMI and AliMeeting. These results suggest that WavLM can provide more powerful representations for the diarization task. Additionally, our model can be further improved by updating WavLM simultaneously with the Conformer training. Compared to the Pyannote3 baseline, when considering no forgiveness collar, our best model achieves relative DER reductions on AMI, AISHELL-4, and AliMeeting of 27.7%, 16.1%, and 32.5%, respectively.
@@ III-C2 Effects of data quantity
Detailed results can be found in Figure 2, where we show the macro-averaged performance across the three datasets when using Fbank, WavLM-frozen, and WavLM-updated features for different training data sizes. As expected, the WavLM features have much better performance than Fbank under each condition. When reducing the available training data, all approaches tend to degrade. However, the effect is more attenuated for WavLM-based systems leading to better results overall. Such a phenomenon indicates that the SSL-based approach is more robust to data scarcity. For an extreme case where only 5% of data is available, the relative DER degradation exceeds 60% when using Fbank features, while for the WavLM-based systems, it is less than 30%. This puts into perspective how much effect the pre-trained SSL model can have on diarization tasks.

### v3

@@ III-B2 Inference
Pyannote3 (baseline) SincNet 21.1 13.9 22.8 13.7 7.7 13.6
Proposed Fbank 19.7 12.5 21.0 12.9 6.9 12.6
WavLM-frozen 17.0 11.7 19.9 10.9 6.1 12.0
WavLM-updated 15.4 11.7 17.6 9.8 5.9 10.2
@@ III-C1 Overall comparison
It is clear that our EEND part has better performance than the Pyannote baseline. Even with the traditional filterbank features, our model significantly outperforms the Pyannote model across the three datasets, suggesting that the Conformer is a better choice for the end-to-end module. By utilizing WavLM, better performance can be obtained, especially for AMI and AliMeeting. These results suggest that WavLM can provide more powerful representations for the diarization task. Additionally, our model can be further improved by updating WavLM simultaneously with the Conformer training. Compared to the Pyannote3 baseline, when considering no forgiveness collar, our best model achieves relative DER reductions on AMI, AISHELL-4, and AliMeeting of 26.6%, 15.8%, and 22.8%, respectively.
@@ III-C2 Effects of data quantity
Detailed results can be found in Figure 2, where we show the macro-averaged performance across the three datasets when using Fbank, WavLM-frozen, and WavLM-updated features for different training data sizes. As expected, the WavLM features have much better performance than Fbank under each condition. When reducing the available training data, all approaches tend to degrade. However, the effect is more attenuated for WavLM-based systems leading to better results overall. Such a phenomenon indicates that the SSL-based approach is more robust to data scarcity. For an extreme case where only 5% of data is available, the relative DER degradation is about 60% when using Fbank features, while for the WavLM-based systems, it is less than 30%. This puts into perspective how much effect the pre-trained SSL model can have on diarization tasks.